A cake recipe calls for 3.75 g of NaHCO3. How many L of CO2 in L is produced in your cake.
1
Expert's answer
2017-03-03T10:27:05-0500
NaHCO3 = 84 g/mol

n (NaHCO3) = 3.75 / 84 = 0.045 mol

NaHCO3 + H2O → H2CO3 + NaOH H2CO3 → H2O + CO2

so 0.045 mol CO2

PV = nRT

1 (V (CO2) = 0.045 · 0.8206 · 463.7

V (CO2) = 17.12 L